Please help me with this word problem:
Use the compound interest formula A=P(1 + r/n)^nt and A=Pe^rt to solve
Find the accumulated value of an investment of $18,000 at 6% compounded semiannually for 11 years
Thank you for your assistance in advance

Answer by mananth(16946)

A=P(1 + r/n)^nt and A=Pe^rt to solve
Find the accumulated value of an investment of $18,000 at 6% compounded semiannually for 11 years
A= 18000(1+(0.06/2))^(11*22)
A = 34,489.86
...
m.ananth@hotmail.ca